When 9.0 g of hydrogen gas, H2, reacts with oxygen gas, O2, 73.0 g of water is produced.    2H2 + O2 --> 2H2O a) What is the theoretical yield of water? b) What is the percentage yield?


2. Powdered zinc metal reacts with sulphur (S8) when heated to produce zinc sulphide.

a) What mass of product can be produced with 25 g of zinc?

b) What is the actual yield if the percentage yield was 80%?


3. The thermite reaction has been used to weld railroad rails, make bombs, and ignite solid rocket fuel. The equation of the reaction is:

a) What is the theoretical yield of iron if 100.0 g of Fe2O3 is used?

b) What is the percentage yield if 52.3 g of iron is actually produced?


1
Expert's answer
2022-01-14T05:28:02-0500

The equation for this reaction is

2H2 + O2 = 2H2O;

Having a mass of hydrogen we can find the quantity substance

n(H2) = m(H2)/M(H2);

M(H2) = 1 *2 = 2 g/mol

n(H2) = 9 g / 2 g/mol = 4.5 moles;

The amount of substance of hydrogen with the amount of substance of water as

a 2 to 2 or 1 to 1.

so n(H2O) = 4.5 moles

Theoretical mass of water will be

n(H2O) = m(H2O)/M(H2O);

M(H2O) = 1 *2 + 16 = 18 g/mol

m (H2O) 4.5 moles * 18 g/mol = 81 g.

so the theoretical yield of water is 81 g.

The percentage yield will be

m(H2O)practical / m(H2O)theoretical = 73.0 g / 81.0 g = 0.9012 or 90.12 %